Calculate the Pressure of 2x10-3 mol gas contained in a 10 ml container at 30 oC. Use the gas constant, R = 0.08214 L atm mol-1 K-1. Keep the answer in atm units.
Answer #1

Universal gas constant = R = 0.08214 L.atm/(mol.K)

Number of moles of the gas = n = 2 x 10-3 mol

Pressure of the gas = P

Volume of the gas = V = 10 ml = 10 x 10-3 L = 0.01 L

Temperature of the gas = T = 30 oC = 30 + 273 K = 303 K

By Ideal Gas Law,

PV = nRT

P(0.01) = (2x10-3)(0.08214)(303)

P = 4.98 atm

Pressure of the gas = 4.98 atm